# An accuracy-runtime trade-off comparison of scalable Gaussian process approximations for spatial data

### v1

@@ B.3 Prediction
B.4 Comparison to exact calculations
@@
Appendix C Simulated data with $N=10^{\prime}000$ and practical range of $0.5$
@@ C.3 Prediction
C.4 Comparison to exact calculations
@@
Appendix D Simulated data with $N=10^{\prime}000$ and practical range of $0.05$
@@ D.3 Prediction
D.4 Comparison to exact calculations
@@
Appendix E Simulated data with $N=100^{\prime}000$ and practical range of $0.2$
E.1 Log-likelihood evaluation
@@ E.3 Prediction
Appendix F Simulated data with $N=100^{\prime}000$ and practical range of $0.5$
@@ F.3 Prediction
Appendix G Simulated data with $N=100^{\prime}000$ and practical range of $0.05$
@@ G.3 Prediction
Appendix H House price dataset
H.1 Prediction
@@ 5 Value 6955 9166 6960 6944 7031
H.3 Comparison to exact calculations
@@
Appendix I Laegern dataset
I.1 Prediction
@@ 5 Value -129805 -107711 -81065 -109549 -125789
Appendix J MODIS 2016 dataset
J.1 Prediction
@@ 5 Value 119174 185366 167307 143126 886374
Appendix K MODIS 2023 dataset
K.1 Prediction
@@ 5 Value 445698 800784 778644 760559
Appendix L Simulated data with $N=100^{\prime}000$ and practical range of $0.05$, single-core setting
@@ 5 Value 129844 142297 139021 136227 131453
Appendix M Simulated data with $N=100^{\prime}000$ and practical range of $0.2$, single-core setting
@@ 5 Value 120195 140101 120356 121315 120767
Appendix N Simulated data with $N=100^{\prime}000$ and practical range of $0.5$, single-core setting
@@ 5 Value 118250 139923 118224 118273 118436
Appendix O House price dataset, single-core setting
@@ 5 Value 6955 9166 6960 6944 7031
O.3 Comparison to exact calculations

### v2

@@ B.1.3 Prediction
B.1.4 Comparison to exact calculations
@@
B.2 Simulated data with $N=10^{\prime}000$ and practical range of $0.5$
@@ B.2.3 Prediction
B.2.4 Comparison to exact calculations
@@
B.3 Simulated data with $N=10^{\prime}000$ and practical range of $0.05$
@@ B.3.3 Prediction
B.3.4 Comparison to exact calculations
@@
B.4 Simulated data with $N=100^{\prime}000$ and practical range of $0.2$
B.4.1 Log-likelihood evaluation
@@ B.4.3 Prediction
B.5 Simulated data with $N=100^{\prime}000$ and practical range of $0.5$
@@ B.5.3 Prediction
B.6 Simulated data with $N=100^{\prime}000$ and practical range of $0.05$
@@ B.6.3 Prediction
B.7 House price dataset
B.7.1 Prediction
@@ 5 Value 6955 9166 6960 6944 7031
B.7.3 Comparison to exact calculations
@@
B.8 Laegern dataset
B.8.1 Prediction
@@ 5 Value -129805 -107711 -81065 -109549 -125789
B.9 MODIS 2016 dataset
B.9.1 Prediction
@@ 5 Value 119174 185366 167307 143126 886374
B.10 MODIS 2023 dataset
B.10.1 Prediction
@@ 5 Value 445698 800784 778644 760559
B.11 Simulated data with $N=100^{\prime}000$ and practical range of $0.05$, single-core setting
@@ B.11.3 Prediction
B.12 Simulated data with $N=100^{\prime}000$ and practical range of $0.2$, single-core setting
@@ B.12.3 Prediction
B.13 Simulated data with $N=100^{\prime}000$ and practical range of $0.5$, single-core setting
@@ B.13.3 Prediction
B.14 House price dataset, single-core setting
@@ 5 Value 6955 9166 6960 6944 7031
B.14.3 Comparison to exact calculations